if (a+b)/a = 6 and (b+c)/c = 9, compute the numerical value of a/c
Answer by venugopalramana(3286)

(a+b)/a = 6 ...
(a/a)+(b/a)=6...
1+b/a=6
b/a=6-1=5..................I.
and (b+c)/c = 9
(b/c)+(/c)=9
b/c=9-1=8......................II
dividing II with I,we get
(b/c)/(b/a)=8/5
(b*a)/(b*c)=8/5
a/c=8/5